Solve the following equation
9y+5 5y+5
----- - -------- = -3
7 2


Answer by checkley77(12844)

(9y+5)/7-(5y+5)/2=-3
[2(9y+5)-7(5y+5)]/14=-3
18y+10-35y-35=-42
-17y=-42+35-10
-17y=-17
y=-17/-17
y=1 answer.